# Automatic AI Model Selection for Wireless Systems: Online Learning via Digital Twinning

### v1

@@ I-A Context, Motivation, and Overview
Figure 2: Wireless resource allocation using AMS: Given the context $c$ defining a network topology, the AMS mapping $\phi=g(c)$ returns a GNN model $\phi$. The GNN model takes the network CSI as input, and it outputs the transmission powers for all the nodes.
@@ I-B Main Contributions
This paper proposes a novel DT-powered online calibration scheme for the optimization of an AMS mapping. As illustrated in Fig. 2, the AMS mapping implements a zero-shot adaptation of the parameters $\phi$ of an AI model on the basis of a context variable $c$ that describes the task of interest. For example, as in Fig. 2, the context information $c$ may encode the current interference graph of the wireless network, and the AI model may be a GNN that determines a power allocation policy.
@@ II Automatic AI Model Selection
In this section, we first define the objective of automatic AI model selection (AMS), and then we review a baseline algorithm for the optimization of AMS that uses only data collected at the physical system.
@@ II-A Automatic Model Selection
While our approach can be applied more broadly, we will use throughout the example illustrated in Fig. 2 for concreteness. In it, the context $c$ represents the topology of a wireless network, and the AI model to be selected is a graph neural network (GNN). The GNN takes channel state information (CSI) of the wireless network as an input, and it outputs the transmission powers for the nodes in the network. The works [30, 31] assumed a GNN model that is designed to perform well across a range of typologies, while the papers[29, 32] demonstrated the performance advantage accrued by adapting the model using data from the current topology via meta-learning. In this work, we aim at automatically mapping context $c$, i.e., topology, to GNN model $\phi$ using the AMS function (1) without requiring data from the current topology. In this sense, AMS implements a form of zero-shot meta-learning[34].
Given a context $c$ at each time step, the performance of an AI model $\phi$ depends on a random vector $x\sim p(x|c)$, whose unknown distribution $p(x|c)$ is context-dependent. The vector $x$ includes the inputs to the AI model, as well as possibly other variables such as feedback signals. For example, for the setting in Fig. 2, the vector $x$ is a realization of the CSI across all the links in the network. We write as $\ell(x|\phi)$ the loss accrued by model $\phi$ for instance $x$, so that the average loss is given by
@@ II-B Calibrating the AMS Mapping
Calibration of the AMS mapping $g(\cdot|\theta)$ aims at producing a parameter vector $\theta$ such that the model $\phi=g(c|\theta)$ yields a small value for the per-context loss in (2). More precisely, writing as $p(c)$ the distribution of the context $c$, we are interested in minimizing the expected value of the per-context loss $\mathcal{L}(\phi|c)$ over the random context $c\sim p(c)$ as
@@
During the calibration phase, at each discrete time step $t=1,2,...$, the system is faced with a context $c_{t}\sim p(c)$, drawn in an independent and identically distributed (i.i.d.) manner over index $t$. By testing the current AMS model parameter $\theta_{t}$, the system estimates the PT loss $L^{{\text{PT}}}(\theta_{t})$ via the empirical average
@@ III-A Digital Twin
The DT is a simulator that, given a context variable $c$, can generate data at some fidelity level $f\in\{1,\dots,F^{\rm{max}}\}$. At fidelity level $f$, the synthetic data $\tilde{x}$ generated by the DT follows the distribution $p(\cdot|c,f)$, which is an approximation of the true context-dependent distribution $p(\cdot|c)$. The quality of the approximation $p(\tilde{x}|c,f)$ improves as the fidelity $f$ increases, but a larger fidelity $f$ entails a larger simulation cost $S_{f}$[38]. Accordingly, the simulation costs are ordered as
@@
At each time step $t=1,2,\dots$ of the calibration phase, the DT simulates the system for a number $M^{\text{DT}}$ context variables $\{c_{m,t}\}_{m=1}^{M^{\text{DT}}}$. Specifically, given a fidelity level $f$, the $N^{\text{DT}}$ data points generated at each context $c_{m,t}$ are distributed as
@@
Using synthetic data, the DT can estimate the PT loss $L^{\text{DT}}(\theta_{t})$ at the current AMS parameter vector $\theta_{t}$ as
@@
If the context variables $\{c_{m,t}\}_{m=1}^{M^{\text{DT}}}$ are drawn i.i.d. from the true context distribution $p(c)$, the average over the DT estimate (11) is
@@
Note that the assumption that the DT is aware of the context distribution $p(c)$ is well justified in settings, such as in Fig. 2, in which the context distribution reflects macroscopic long-term trends such as the spatial density of devices [30, 31, 33]. Importantly, due to the discrepancy between DT distribution $p(\cdot|c,f)$ and the real-world distribution $p(\cdot|c)$, the expected DT loss (13) is generally different from the PT loss $L^{{\text{PT}}}(\theta)$ in (4), i.e.,
@@ III-B Naïve DT-Powered AMS
The simplest way for the DT to use both synthetic and real-world data is to weigh both sources of data equally, that is, by treating the synthetic data as if it was genuine real-world data. Using the empirical losses (5) and (11), this yields to the estimate
@@ III-C DT-Powered AMS
The key idea is that the DT uses data simulated for the current context $c_{t}$ in order to evaluate the bias caused by simulation errors. To this end, at each time step $t$ of the calibration phase, the PT informs the DT about the current context variable $c_{t}$. With this information, the DT simulates data $\{\tilde{x}_{n,t}\}_{n=1}^{N^{\text{DT}}}$ with $\tilde{x}_{n,t}\mathop{\sim}\limits^{\text{i.i.d.}}p(\tilde{x}|c_{t},f)$ under the current context $c_{t}$ for the selected fidelity level $f$. If the simulation budget in (10) allows, the $M^{\text{DT}}-1$ remaining contexts $\{c_{m,t}\}_{m=1}^{M^{\text{DT}}-1}$ are chosen i.i.d. from distribution $p(c)$ as in N-DT-AMS. Using the synthetic data $\{\tilde{x}_{n,t}\}_{n=1}^{N^{\text{DT}}}$ generated for the current context $c_{t}$, the DT estimates the average loss as
@@ IV-B Adapting the Hyperparameters
where $V^{\text{PT}}(\theta_{t})$, $V^{\text{DT}}(\theta_{t})$, and $V^{\text{DT}\rightarrow\text{PT}}(\theta_{t})$ are defined as the variance of the estimates $\hat{L}^{\text{PT}}(\theta_{t})$ in (5), $\hat{L}^{\text{DT}}(\theta_{t})$ in (20), and $\hat{L}^{\text{DT}\rightarrow\text{PT}}(\theta_{t})$ in (17), respectively, i.e.,
$V^{\text{PT}}(\theta_{t})$ $=\mathbb{E}[(\hat{L}^{\text{PT}}(\theta_{t})-L^{\text{PT}}(\theta_{t}))^{2}],$ (28)
@@ V-A1 Scenario
As depicted in Fig. 5, the physical twin consists of a network with $K$ single-antenna transceiver pairs $\{(T_{t}^{k},R_{t}^{k})\}^{K}_{k=1}$ located in a square deployment area with sides of length $A$. We denote as $T_{t}^{k}$ and $R_{t}^{k}$ the positions of the $k$-th transmitter and receiver, respectively, at each time step $t$ of the calibration phase. At time step $t$, the position $T_{t}^{k}$ of the $k$-th transmitter is randomly and uniformly selected in the deployment area, while the corresponding receiver’s position $R_{t}^{k}$ is randomly and uniformly chosen within a two-dimensional torus centered at $T_{t}^{k}$ with radii $d_{\rm{min}}$ and $d_{\rm{max}}$.
@@
where $L(d^{j,k}_{t})$ is the path-loss at distance $d^{j,k}_{t}$ [36]; $\eta$ is the transmit antenna power gain; $\sigma_{s}^{2}$ is the shadowing power; $\alpha_{j,k}\sim\mathcal{U}[-\pi,\pi]$ is the random phase of the line-of-sight (LoS) component; $\hat{h}_{j,k}\sim\mathcal{CN}(0,1)$ models Rayleigh fading for the non line-of-sight (NLoS) component; and $\kappa$ is the Rician factor determining the strength of the LoS component as compared to the NLoS component. We denote as $\mathbf{H}_{t}\sim p(\mathbf{H}|\mathbf{D})$ the overall CSI matrix which has $(j,k)$-th element given by (34), i.e., $[\mathbf{H}_{t}]_{j,k}=h^{j,k}_{t}$, with $p(\mathbf{H}|\mathbf{D})$ being the conditional distribution over the CSI matrix given the distance matrix $\mathbf{D}$.
@@ V-A2 Data Generation
At each time step $t$, the system collects $N^{\text{PT}}$ CSI matrices $\{\mathbf{H}_{n,t}\}_{n=1}^{N^{\text{PT}}}\mathop{\sim}\limits^{\text{i.i.d.}}\\ p(\mathbf{H}|\mathbf{D}_{t})$ given the current distance matrix $\mathbf{D}_{t}\sim p(\mathbf{D})$. This corresponds to the data points $\{x_{n,t}\}_{n=1}^{N^{\text{PT}}}$ described in the text. Having received information about the context $\mathbf{D}_{t}$, the DT generates data for context $\mathbf{D}_{t}$ at a fidelity level dictated by the degree to which the topology modelled by the simulator. Specifically, given a fidelity level $f\in(0,1)$, the DT simulates a topology characterized by a mismatched distance matrix containing a fraction $f$ of the links as compared to the original distance matrix $\mathbf{D}$. To model this real-to-sim discrepancy, each link of the $K(K-1)$ links is dropped independently with probability $1-f$. The resulting simulation cost $S_{f}$ is quantified as
@@
TABLE I: Correspondence between Variables in Main Text and Experiment
@@
In summary, the corresponding relationship between the variables in the main text and the simulation under the study is summarized in Table I. Specific system parameters in our simulation scenario and the data generation parameters are summarized in Table II and Table III, respectively.
TABLE II: System Parameters
@@
TABLE III: Data Generation Parameters at Time Step $t$
@@ V-B AMS Mapping and Graph Neural Network-Based Power Control
Given a context provided by the distance matrix $\mathbf{D}_{t}$, the AMS mapping aims at determining an AI app that can maximize the sum-rate of the $K$ users by controlling the transmission powers as a function of the CSI. To this end, the AMS function maps distance matrix $\mathbf{D}$ to the parameters $\phi$ of the power-controlling GNN. The GNN follows the wireless communication graph convolution network (WCGCN) architecture in [30]. The GNN takes as input the $K\times K$ CSI matrix $\mathbf{H}$ to output a $K\times 1$ transmit power vector $P$. The loss function in (2) is specified as the negative sum-rate, i.e.,
@@
In Algorithm 1, we set $\lambda_{0}=1$ and $\mu_{0}=0.5$, and the window size $W$ is decreased exponentially over time $t$ from $40$ to $5$, with the window size decreased every time the performance does not improve over two consecutive time steps. The choice of initial hyperparameters $\lambda_{0}\textgreater\mu_{0}$ imposes a higher reliance to the low-variance DT-estimated loss (20) at the initial training stage, which we found empirically to be useful for optimization stability, in a manner similar to the observation reported in [40]. The learning rate in (7), (16), (22), and (24) are all set as $0.015$.
@@ VI Simulation Results
Finally, to set a performance upper bound, we consider the weighted minimum mean-square error (WMMSE) power control algorithm [43], a benchmark model-based power control scheme that is optimized separately for each CSI matrix $\mathbf{H}$. For each scheme, we report the mean performance and $95\%$ confidence intervals obtained averaging the results over $5$ random simulation seeds. Simulation code is available at https://github.com/qiushuo0913/DT-powered-AMS.

### v2

@@ I-B Main Contributions
This paper proposes a novel DT-powered online calibration scheme for the optimization of an AMS mapping. As illustrated in Fig. 2, the AMS mapping implements a zero-shot adaptation of the parameters $\phi$ of an AI model on the basis of a context variable $c$ that describes the task of interest. For example, the context information $c$ may encode the matrix of geographic distances obtained from the locations of the devices, and the AI model may be a GNN that determines a power allocation policy based on channel state information (CSI).
@@
Figure 2: Wireless resource allocation using AMS: Given the context $c$ describing the matrix of geographical distances among the network nodes, i.e., the network topology, the AMS mapping $\phi=g(c)$ returns a GNN model $\phi$. The GNN model takes the current network CSI as input, and it outputs the transmission powers for all the nodes.
@@ II Automatic AI Model Selection
In this section, we first define the objective of AMS, and then we review a baseline algorithm for the optimization of AMS that uses only data collected at the physical system. We summarize the main notations used throughout this paper in Table I.
TABLE I: Notation Used in This Paper
Symbol Definition
$c$ context variable
$c_{t}$ context variable collected from PT at time step $t$
$c_{m,t}$ $m$-th context variable generated by DT at time step $t$
$\phi$ parameter vector of the AI model
$\theta$ parameter vector of the AMS mapping
$x$ data
$\ell(x|\phi)$ loss accrued by model $\phi$ for data $x$
$g(\cdot|\theta)$ AMS mapping parameterized by a vector $\theta$
$x_{n,t}$ $n$-th data collected from PT at time step $t$
$\tilde{x}_{n,t}$ $n$-th data generated by DT at time step $t$
$\tilde{x}_{m,n,t}$ $n$-th data generated by DT based on context variable $c_{m,t}$
$N^{\text{PT}}$ number of data points collected from PT at each time step
$N^{\text{DT}}$ number of data points generated by DT at each time step
$M^{\text{DT}}$ number of context variables generated by DT at each time step
$f$ fidelity level of DT
$S_{f}$ simulation cost of DT at fidelity level $f$
$S^{\text{DT}}$ total simulation budget at DT
$p(c)$ distribution of context variable $c$
$p(\cdot|c)$ data generating distribution at PT given context $c$
$p(\cdot|c,f)$ simulated data-generating distribution $p(\cdot|c)$ at DT under fidelity given context $c$
@@ II-A Automatic Model Selection
While our approach can be applied more broadly, we will use throughout the example illustrated in Fig. 2 for concreteness. In it, the context $c$ represents the topology of a wireless network, and the AI model to be selected is a graph neural network (GNN). The GNN takes CSI of the wireless network as an input, and it outputs the transmission powers for the nodes in the network. The works [30, 31] assumed a GNN model that is designed to perform well across a range of typologies, while the papers[29, 32] demonstrated the performance advantage accrued by adapting the model using data from the current topology via meta-learning. In this work, we aim at automatically mapping context $c$, i.e., topology, to GNN model $\phi$ using the AMS function (1) without requiring data from the current topology. In this sense, AMS implements a form of zero-shot meta-learning[34].
Given a context $c$ at each time step, the performance of an AI model $\phi$ depends on a random vector $x\sim p(x|c)$, whose unknown distribution $p(x|c)$ is context-dependent. The vector $x$ includes the inputs to the AI model, as well as possibly other variables such as feedback signals. For example, for the setting in Fig. 2, the vector $x$ is a realization of the CSI across all the links in the network. The distribution $p(x|c)$ is practically unknown, and it can be only approximated via modeling strategies, including ray tracing in the case of CSI[47]. We write as $\ell(x|\phi)$ the loss accrued by model $\phi$ for instance $x$, so that the average loss is given by
@@ II-B Calibrating the AMS Mapping
Calibration of the AMS mapping $g(\cdot|\theta)$ aims at producing a parameter vector $\theta$ such that the model $\phi=g(c|\theta)$ yields a small value for the per-context loss in (2). More precisely, writing as $p(c)$ the marginal distribution of the context variable $c$, we are interested in minimizing the expected value of the per-context loss $\mathcal{L}(\phi|c)$ over the random context $c\sim p(c)$ as
@@
During the calibration phase, at each discrete time step $t=1,2,...$, the system is faced with a context $c_{t}\sim p(c)$. Note that the context variables $c_{t}$ may be correlated across the time index $t=1,2,\dots$, as long as they share the same marginal distribution $p(c)$. For example, the sequence of context variables $c_{1},c_{2},\dots$, may follow a Markov model with stationary distribution $p(c)$. By testing the current AMS model parameter $\theta_{t}$, the system estimates the PT loss $L^{{\text{PT}}}(\theta_{t})$ via the empirical average
@@ III-A Digital Twin
Following the conventional definition of digital twining[46], in this paper, the DT is a simulator of the PT (see also [11, 13, 16, 18]). Specifically, given a context variable $c$, the DT can generate synthetic data at some fidelity level $f\in\{1,\dots,F_{\rm{max}}\}$. At fidelity level $f$, the synthetic data $\tilde{x}$ generated by the DT follows the distribution $p(\cdot|c,f)$, which is an approximation of the true context-dependent distribution $p(\cdot|c)$. The quality of the approximation $p(\tilde{x}|c,f)$ improves as the fidelity $f$ increases, but a larger fidelity $f$ entails a larger simulation cost $S_{f}$[38]. Accordingly, the simulation costs are ordered as
@@
At each time step $t=1,2,\dots$ of the calibration phase, the DT simulates the system for a number $M^{\text{DT}}$ context variables $\{c_{m,t}\}_{m=1}^{M^{\text{DT}}}$ following the marginal distribution $p(c)$. Specifically, given a fidelity level $f$, the $N^{\text{DT}}$ data points generated at each context $c_{m,t}$ are distributed as
@@
Using synthetic data, the DT can estimate the DT loss $L^{\text{DT}}(\theta_{t})$ at the current AMS parameter vector $\theta_{t}$ as
@@
As long as the context variables follow the marginal, stationary, distribution $p(c)$, the average over the DT estimate (11) is
@@
Note that the assumption that the DT is aware of the stationary context distribution $p(c)$ is well justified in settings, such as in Fig. 2, in which the distribution $p(c)$ reflects macroscopic long-term trends, including the spatial density of devices [30, 31, 33]. That said, as further discussed in the next section, it is also possible to account for differences between the true distribution $p(c)$ and the sampling distribution assumed by the DT. Furthermore, the contexts $\{c_{m,t}\}_{m=1}^{M^{\text{DT}}}$ generated by the DT may potentially be correlated in order to capture the memory of the real context process. This is elaborated on in Section VI-D.
Importantly, due to the discrepancy between DT distribution $p(\cdot|c,f)$ and the real-world distribution $p(\cdot|c)$, the expected DT loss (13) is generally different from the PT loss $L^{{\text{PT}}}(\theta)$ in (4), i.e.,
@@ III-B Naïve DT-Powered AMS
The simplest way for the DT to use both synthetic and real-world data is to weigh both sources of data equally, that is, by treating the synthetic data as if it had been genuine real-world data. Despite its simplicity, this approach plays an important role in semi-supervised learning literature as a strong baseline (see, e.g., [40]).
Using the empirical losses (5) and (11), this yields to the estimate
@@ III-C DT-Powered AMS
The key idea is that the DT uses data simulated for the current context $c_{t}$ in order to evaluate the bias caused by simulation errors. To this end, at each time step $t$ of the calibration phase, the PT informs the DT about the current context variable $c_{t}$. With this information, the DT simulates data $\{\tilde{x}_{n,t}\}_{n=1}^{N^{\text{DT}}}$ with $\tilde{x}_{n,t}\mathop{\sim}\limits^{\text{i.i.d.}}p(\tilde{x}|c_{t},f)$ under the current context $c_{t}$ for the selected fidelity level $f$. If the simulation budget in (10) allows, the $M^{\text{DT}}-1$ remaining contexts $\{c_{m,t}\}_{m=1}^{M^{\text{DT}}-1}$ are chosen from the marginal distribution $p(c)$ as in N-DT-AMS. Using the synthetic data $\{\tilde{x}_{n,t}\}_{n=1}^{N^{\text{DT}}}$ generated for the current context $c_{t}$, the DT estimates the average loss as
@@ IV-B Adapting the Hyperparameters
where $V^{\text{DT}}(\theta_{t})$, and $V^{\text{DT}\rightarrow\text{PT}}(\theta_{t})$ are defined as the variance of the estimates $\hat{L}^{\text{DT}}(\theta_{t})$ in (20), and $\hat{L}^{\text{DT}\rightarrow\text{PT}}(\theta_{t})$ in (17), respectively, i.e.,
@@
The optimal values (27) rely on knowledge at the DT about the marginal context distribution $p(c)$. As shown in Appendix A, it is possible to generalize the optimal solution (27) to a setting in which the DT only has an available estimate $\hat{p}(c)$ of the true context distribution $p(c)$.
@@ V-A1 Scenario
As depicted in Fig. 5, following a common formulation of the power control problem in the presence of multiple transceiver pairs[30, 31], the PT consists of a network with $K$ single-antenna transceiver pairs $\{(T_{t}^{k},R_{t}^{k})\}^{K}_{k=1}$ located in a square deployment area with sides of length $A$. We denote as $T_{t}^{k}$ and $R_{t}^{k}$ the positions of the $k$-th transmitter and receiver, respectively, at each time step $t$ of the calibration phase. At time step $t$, the position $T_{t}^{k}$ of the $k$-th transmitter is randomly and uniformly selected in the deployment area, while the corresponding receiver’s position $R_{t}^{k}$ is randomly and uniformly chosen within a two-dimensional torus centered at $T_{t}^{k}$ with radii $d_{\rm{min}}$ and $d_{\rm{max}}$.
@@
where $L(d^{j,k}_{t})$ is the path-loss at distance $d^{j,k}_{t}$ [36]; $\eta$ is the transmit antenna power gain; $\sigma_{s}^{2}$ is the shadowing power; $\alpha_{j,k}\sim\mathcal{U}[-\pi,\pi]$ is the random phase of the line-of-sight (LoS) component; $\hat{h}_{j,k}\sim\mathcal{CN}(0,1)$ models Rayleigh fading for the non line-of-sight (NLoS) component; and $\kappa$ is the Rician factor determining the strength of the LoS component as compared to the NLoS component. We denote as $\mathbf{H}_{t}\sim p(\mathbf{H}|\mathbf{D})$ the overall CSI matrix which has the $(j,k)$-th element given by (33), i.e., $[\mathbf{H}_{t}]_{j,k}=h^{j,k}_{t}$, with $p(\mathbf{H}|\mathbf{D})$ being the conditional distribution over the CSI matrix given the distance matrix $\mathbf{D}$.
@@ V-A2 Data Generation
At each time step $t$, the system collects $N^{\text{PT}}$ CSI matrices
$\{\mathbf{H}_{n,t}\}_{n=1}^{N^{\text{PT}}}\mathop{\sim}\limits^{\text{i.i.d.}}p(\mathbf{H}|\mathbf{D}_{t})$ given the current distance matrix $\mathbf{D}_{t}\sim p(\mathbf{D})$. This corresponds to the data points $\{x_{n,t}\}_{n=1}^{N^{\text{PT}}}$ described in the text. Having received information about the context $\mathbf{D}_{t}$, the DT generates data for context $\mathbf{D}_{t}$ at a fidelity level dictated by the degree to which the topology modelled by the simulator. Specifically, given a fidelity level $f\in(0,1)$, the DT simulates a topology characterized by a mismatched distance matrix containing a fraction $f$ of the links as compared to the original distance matrix $\mathbf{D}$. To model this real-to-sim discrepancy, each link of the $K(K-1)$ links is dropped independently with probability $1-f$. The resulting simulation cost $S_{f}$ is quantified as
@@
TABLE II: Correspondence between Variables in Main Text and Experiment
@@
In summary, the corresponding relationship between the variables in the main text and the simulation under the study is summarized in Table II. Specific system parameters in our simulation scenario and the data generation parameters are summarized in Table III and Table IV, respectively.
TABLE III: System Parameters
@@
TABLE IV: Data Generation Parameters at Time Step $t$
@@ V-B AMS Mapping and Graph Neural Network-Based Power Control
Given a context provided by the distance matrix $\mathbf{D}_{t}$, the AMS mapping aims at determining an AI app that can maximize the sum-rate of all $K$ users by controlling the transmission powers as a function of the CSI. To this end, the AMS function maps distance matrix $\mathbf{D}$ to the parameters $\phi$ of the power-controlling GNN. The GNN follows the wireless communication graph convolution network (WCGCN) architecture in [30]. The GNN takes as input the $K\times K$ CSI matrix $\mathbf{H}$ to output a $K\times 1$ transmit power vector $P$. The loss function in (2) is specified as the negative sum-rate as in [29, 30, 31], i.e.,
@@
In Algorithm 1, we set $\lambda_{0}=1$ and $\mu_{0}=0.5$, and the window size $W$ is decreased exponentially over time $t$ from $40$ to $5$, with the window size decreased every time the performance does not improve over two consecutive time steps. The choice of initial hyperparameters $\lambda_{0}\textgreater\mu_{0}$ imposes a higher reliance to the low-variance DT-estimated loss (20) at the initial training stage, which we found empirically to be useful for optimization stability, in a manner similar to the observation reported in [40]. The learning rate in (7), (16), (22), and (24) are all set as $0.015$. We refer to Appendix C for details on training.
@@ VI Simulation Results
Finally, to set a performance upper bound, we consider the weighted minimum mean-square error (WMMSE) power control algorithm [43], a benchmark model-based power control scheme that is optimized separately for each CSI matrix $\mathbf{H}$. For each scheme, we report the mean performance and $95\%$ confidence intervals obtained by averaging the results over $5$ random simulation seeds. Simulation code is available at https://github.com/qiushuo0913/DT-powered-AMS.